Q\Determine the following ODEs are exact or not? Why?
(2xy)dx+2y^2dy=0
1
Expert's answer
2020-10-26T20:04:21-0400

Differential equation "Mdx+Ndy=0" is exact if and only if "\frac{\partial M }{\partial y}=\frac{\partial N }{\partial x}".

"\frac{\partial M }{\partial y}=2x"

"\frac{\partial N }{\partial x}=2y^2."

ODE is not exact.